cerna.gif (1167 bytes)

List of Actions from MPP Meetings

lhc_logo.gif (4175 bytes)


MPP

Description

Responsible

Comments

149th (23.06.2017)
  • Make a final proposal for the implementation of the RQS in the maskable interlocks during TS#2 via an ECR.

MPE-MI

Was presented back in MPP#155, RQS circuits were kept in the NO_IMPACT category.
149th (23.06.2017)
  • Determine how many times the RQS circuits went into fault in the past year.

M. Valette

The logged data on the RQS circuits suggests that the RQS would lead to 4 dumps a year on average. Which would have a visible effect on availability.
146th (12.05.2017)
  • Prepare a detailed procedure for the ADT AC-dipole test and provide sample data to prepare for the online post-processing.

B. Lindstrom, D. Valuch

The test was performed successfully on 21.05.2017.
146th (12.05.2017)
  • Verify with Rogelio and the tune spread measurements team what intensities and ADT settings they require for their measurements.

MPP

Was tested in an MD and is now implemented in the tune measurement tool.
145th (21.04.2017)
  • Provide waveforms to COLL from MKDf.B2 erratic (to compare with asynch beam dump type 2).

N. Magnin, R. Bruce

The waveforms were provided by N. Magnin. As the event is over a year old and did not occur again it will be studied in the future.
145th (21.04.2017)
  • Schedule a talk in MPP on the status of the Roman Pots, the presentation should include strategy during the intensity ramp up.

S. Jakobsen, M. Deile, M. Trzebinski

Done in MPP#147.
145th (21.04.2017)
  • Organise a joint meeting of the CWG and the MPP at the end of May to finalize the BPM and Roman Pots statuses during the intensity ramp up.

MPP

Done in MPP#147.
145th (21.04.2017)
  • Request proposal from TE-MPE-MI on CIBDS performance during reliability run and sent invitation for after 8:30 meeting.

CIBDS team

Reliability run was successfully performed during commissioning. No CIBDS failures were noticed during the 2016/17 LBDS reliability run.
145th (21.04.2017)
  • Confirm starting configuration for AGK 288b or 144b.

Machine Coordinators

Operation started with 144b.
145th (21.04.2017)
  • Distribute condensed list of MPS tests to be performed with beam.

MPP

Done during commissioning.
144th (07.04.2017)
  • The turn by turn interlock of the SPS BIs in BA6 will be disconnected and the CIBU and BI hardware removed.

BIS Team

The " turn by turn" interlock is now removed from the BIS input 9 in BA6: - CIBU removed from User rack - Cable disconnected from the BIC - Disable jumpers in place on the CIBM - CCD modified Restart the java BIS GUI from the different SPS console to update this latter.
144th (07.04.2017)
  • Prepare an overview list of all threshold changes for 2017.

C. Xu

Was presented in BLMTWG#51.
143rd (31.03.2017)
  • SIS: TDE nitrogen pressure interlock: Change gauge used for dumping in case of too low pressure (< 1 bar) to the newly installed one close to dump volume (as proposed by last MPP); introduce new injection inhibit if either of the two pressure gauges measures a pressure below 1.025 bar. In both cases implement waiting time of 1 min to avoid triggering on communication issues or wrong pressure readings.

BE-OP

The SIS interlock was implemented. It is active in UA65 but deactivated for UA63 following the conclusions from MPP#153 focussiong on operation of the TDE in air.
143rd (31.03.2017)
  • Define interlock logic for DOROS BPMs at collimators (what to do if no data or data inconsistent etc).
  • Define requirements for DOROS BPM post mortem data and provide them to BE-BI

Collimation, BE-OP, MPP

Was presented back in MPP#156 with a follow up action on definition of use cases.
141st (03.03.2017)
  • Verifiy the compatibility of the foreseen commissioning date for the 24bit ADC DCBCT (first week of April)

MPP

Done.
140th (17.02.2017)
  • What is the time required until the TDE bloc can be fully emerged in air due to a worst case leak?
  • What is the maximum temperature allowed in the TDE block, when emerged in air before it is severely damaged?
  • Compare the reliability of a (new) hardware and the existing SIS interlock of the TDE N2 pressure data?
  • G. Pigny, G. Bregliozzi
  • EN-STI
  • A. Apollonio, G. Pigny, G. Breggliozzi
Was the focus of MPP#153. A memo was published on limits for operation of TDE in air.
139th (03.02.2017)
  • Report on the outcome of the discussion on the PM-XPOC data from the BPMS.

Jan

The data was moved to the general XPOC, ideally it is available in PostMortem under the BPMLHC system.
139th (03.02.2017)
  • Finalize the specifications for the dynamic range of the interlock BPMs.

BE-BI, TE-ABT, MPP

Was presented back in MPP#155, with a follow up action.
136th (11.11.2016)
  • Perform a reliability run on the MKB with lower voltage during the EYETS and study the new failure case. Determine how critical for the dump window and absorbers it is.

ABT

Presented in Chamonix'17. Will be presented back in MPP during spring.
136th (11.11.2016)
  • Transmit the proposal for the increased asynchronous beam dump delay to Etienne Carlier and Nicolas Magnin so they can comment on the number.

S. Gabourin

Done.
136th (11.11.2016)
  • Come back before the EYETS with a testing plan for the new CIBDS.

S. Gabourin

Done, came back for MPP141.
135th (28.10.2016)
  • Define, plan and perform reliability run for the CIBDS V5.

ABT, MPE

Done, came back for MPP136.
135th (28.10.2016)
  • Check the timing of the CIBDS retrigger pulse so as to decide if we need to increase the TDU delay time of 270 µs for the CIBDS asynchronous beam dump trigger. Also don’t merge A and B signals for user permits.

S. Gabourin

Done, came back for MPP136.
135th (28.10.2016)
  • Verify the reliability of the wire voltage measurement system (Adriana). Specify connection to WIC.

A. Rossi, J. Uythoven

Done.
134th (30.09.2016)
  • Study possibility to connect new TDE pressure measurement to VAC controller in UA63/UA67 and implement a HW interlock via the vacuum system.

G. Bregliozzi

Done, came back for MPP140.
134th (30.09.2016)
  • Prepare specifications for WIC interlocking of wire temperature via resistance/voltage supervision and present to MPP after the protection electronics and algorithm has been designed and tested.

A. Rossi, Collimation

Done, came back for MPP135.
132nd (26.08.2016)
  • Further PC interlock development is needed: integration of D1 and arc/IP COD families. Clean what is left of PC interlocking in the SIS and integrate it in the PC interlock.

J. Wenninger, K. Fuchsberger

Done in 2015. The D1 was not integrated in the Pc interlock becasue it does not fit the scope at the moment.
133rd (16.09.2016)
  • Derive updated FMCM thresholds from the tracking studies and the power converter tests for after the EYETS.
  • Study / test sensitivity and ripple of new power converters.
  • Study the effect of a RD34 failure on the orbit and the associated FMCM thresholds.

MPE-MI, M. Valette

The new converteers show no perturbation in case of electric glitches as demonstrated in 2017. The FMCM thresholds were kept to initial values but will have to be re-evaluated in case optics with lower beta* are used.
132nd (26.08.2016)
  • Review make-rule strategy for the PC interlock.

MPP

Was presented back in MPP#157, with updated rules for 2018.
129th (10.06.2016)
  • Make a proposal for the maximum excitation windows length and amplitude.

D. Valuch, R. Tomas

Done. MP test during commisssioning to be performed.
129th (10.06.2016)
  • Make a proposal for the test at injection and at flat top, mentioning how many bunches will be excited and with which amplitude.

E. Metral

MD was performed succesfuly.
128th (03.06.2016)
  • Prepare a list with loss maps required after the XRP intervention. From that the machine coordinators can derive a list of required fills.

Collimation, S. Redaelli

Done.
127th (27.05.2016)
  • Propose interlock limits of the Quad PC interlock to MPP.

K. Fuchsberger

Done, came back for MPP132.
127th (27.05.2016)
  • Review experience with Coll BPMs and propose interlock settings and an appropriate moment to activate.
  • Propose local orbit interlock for the TOTEM bump.

Coll WG

Was concluded in MPP#152.
127th (27.05.2016)
  • Check PC interlock server status and experience with stability, and Kajetan. Afterwards activate the PC interlock for the CODs around TS1.

M. Zerlauth

Done, came back for MPP132.
127th (27.05.2016)
  • Prepare procedure for the proposed study:
    • new insertion proposal (to be used after TS1)
    • measure acceptance versus TCL positions

AFP

Under aproval, see EDMs document #1690876.
126th (13.05.2016)
  • Discuss the protection and hierarchy (incl. phase advance between MKD and TCTs) for 2.5km optics.

MPP, Coll, OP

Done.
125th (22.04.2016)
  • Propose and implement a solution to achieve a redundant pressure interlocking system for TDE.

VSC & STI

Done, came back for MPP140.
125th (22.04.2016)
  • Request to the logging service a modification to archive the AFP positions, limits ect in TIMBER.

OP & AFP

The relevant variables were added to TIMBER DB.  
125th (22.04.2016)
  • Work on the TDE pressure alarm proposal to implement it in Big Sister.

VSC & OP

Done.
125th (22.04.2016)
  • Understand where the 1.3 bar upper pressure limit for the TDE was derived from.

ABT

The TDE tank is not a pressure vessel. 
124th (08.04.2016)
  • Prepare settings EDMS document for TOTEM and AFP and verify with Belen.

TOTEM & AFP

Link to EDMs document.  
124th (08.04.2016)
  • Check with Gianluca, to disentangle the three experiments (TOTEM, ALFA, AFP), as for the moment if the FESA of one of these three is down, none can be used in the GUI for movement.

ALFA, TOTEM & AFP

Done, the links in CCM have been updated, there is now a new submenu called Roman Pots within Collimators.  
124th (08.04.2016)
  • Verify the logic for not-installed AFP roman pots after home switch signal forced true.

AFP

Came back in MPP 127, see minutes. 
124th (08.04.2016)
  • Initiate discussion on reviewing upgrade needs for the triggering electronics of the LBDS.

M. Fraser

Came back on MPP124. 
124th (08.04.2016)
  • Perform lab tests on MKD PC stacks with 7 TeV to confirm criticality of type 3 erratic.

M. Fraser

Came back on MPP124. 
123rd (18.03.2016)
  • Contact BIS users without automatic test mode, verify if changes have been performed during the YETS and recommission if necessary

MPP

 
123rd (18.03.2016)
  • Organize review on Linac4 MP in June in preparation of the 160MeV commissioning and half sector test

D. Nisbet

A special half day MPP on Linac4 Machine Protection will be held on July 29th 2016. 
123rd (18.03.2016)
  • Propose the watchdog interlock current loss threshold for the 50MeV phase of the LINAC4 Commissioning

D. Nisbet

A special half day MPP on Linac4 Machine Protection will be held on July 29th 2016. 
121st (19.02.2016)
  • Define the different steps of the FGC lite's reduced hardware commissioning.

S. Uznanski

Was not installed during TS2, will be installed during EYETS, HW commissioning to be discussed then.
121st (19.02.2016)
  • Advance with AFP interlock tests, create some documentation and give a summary on the interlock tests in the 122nd MPP

M. Trzebinski

 
121st (19.02.2016)
  • Trigger expert discussion to come to a machine protection statement, is phase interlocking mandatory to go to 40cm or not?

K. Fuchsberger, MPP

 
121st (19.02.2016);
  • Increase the FGC lite ramp down delay period to 10s to allow the OFB and SIS to take action prior to a potential power cycle

S. Uznanski

Specifications in the EDMs document and firmware updated
120th (15.01.2016)
  • Assess with ABT the possibility of further tightening down the interlock windows for the injection BETS to reduce the possibility of beam impacts on the TDI

E. Carlier

Cancelled. The BETS MKI were not changed.
120th (15.01.2016)
  • Follow-up in collaboration with BI and the BIS team the reliability of intensity input to be possibly included in the interlock logic

V. Kain

Safe Machine Parameters to be reviewed. Presented back in MPP141.
119th (23.10.2015)
  • Finalize tests on BIS fiber optics causing dumps because of attenuation

S. Gabourin

Presented on a later MPP. 
118th (09.10.2015)
  • verify whether a filter is installed on Q5R8 and all DB definitions are coherent

BLM team

 
118th (09.10.2015)
  • verify feasibility to maintain 2 BLECs formware versions for inection blinding for an additional year (up to EYETS 2016/17)

BLM team

Decision by the end of 2016, an LMc note was requested.
117th (25.09.2015);
  • Agree on COLL/RP settings for 90m run

TOTEM, ALFA,  COLL

Done and updated in MP document for RP settings in 2015
117th (25.09.2015)
  • Update on LINAC4 commssioning steps and calidation for 50MeV

D.Nisbet, BIS team

 
116th (04.09.2015);
  • Verify loss signals with independent movements of TCL5 and TCL6

TOTEM

Done, report in future MPP meeting
116th (04.09.2015);
  • Roll back PXI software version to fully approved one

TOTEM

Done by S.Ravat
116th (04.09.2015);
  • Verify whether the magnitude of the vacuum spike varies with intensity

TOTEM

verified by M.Deile and no correlation found
115th (28.08.2015);
  • Verify the scaling for WS-MQY (cell 6) thresholds for scanning with wire scanner of 15b. 

BLM team

Confirmed OK up to 15b at 6.5TeV
114th (21.08.2015);
  • Write small ECR for BCCM activation and propose to LMC. 

MPP

Done, activated early September
113th (14.08.2015);
  • Verify unusally low signals on monitor BLMQI.06L8.B1E30 for eventual NCs in cabling or functionality

BLM team

The verification of the filter installation was done later in 2015. 
112th (31.07.2015);
  • Review of the operational procedure in case of a non-working dump and start of related basic tests before TS2.

MPP

First update done in MPP, 2nd iteration in MPP #119
112th (31.07.2015);
  • Confirm relocation of direct BLMs hardware in a new crate.

E. Effinger

 
112th (31.07.2015);
  • Provide a list of reference people to be contacted for the approval of changes of master thresholds for run 2.

MPP

 
112th (31.07.2015);
  • Finalize documentation for the testing procedure of direct BLMs to the LBDS.

E. B. Holzer

Integarted in BLM comissioning doc
111th (09.07.2015);
  • Define a strategy and a responsible person for the commissioning of the full interlock chain (user logic/thresholds + BIS response).

A. Lombardi, J-B. Lallement

Done, D.Nisbet has been confirmed as MP linkperson.
111th (09.07.2015);
  • Define a person responsible for the Linac4 SIS implementation.

A. Lombardi, J-B. Lallement

Done, D.Nisbet has been confirmed as MP linkperson who will also take care of the SIS implementation.
110th (19.06.2015);
  • Define sequences to set collimators.

S. Redaelli

 
110th (19.06.2015);
  • Finalize documentation for the testing procedure of direct BLMs to the LBDS.

E. B. Holzer

Done and successfully validated with beam.
109th (29.05.2015);
  • Follow-up BCT/SMP performance, comparing signals from the DCBCT with the SBF transitions.

M. Ludwig, S. Gabourin

Solved by implementation of 24bits BCT in SPS and LHC.
109th (29.05.2015);
  • Update on implementation of virtual beta* limits.

A. Masi

Presented by J. Wenninger in the summer.
108th (24.04.2015);
  • Discuss BPM settings for doublet bunches.

T. Lefevre

Proposal made and appllied for first doublet tests.
108th (24.04.2015);
  • Check error found in the BPM IR6 software on the FESA side.

T. Lefevre

 
107th (24.04.2015);
  • Check visibility of XRP PXIs via DIAMON in the CCC.

M. Deile

Done, not accessible.
106th (17.04.2015);
  • Create documentation on the arming sequence of BIS and SMP and the inherent delays to be respected

MPP, D.Calcoen

Done, the documentation can be found there.
106th (17.04.2015);
  • Check timing and responses of PIC, QPS and BLMs for the beam induced quenches observed at the beginning of the 2015 run. In most of the quenches the dump was triggered by only one system. The orbit data was not available up to ~ 14/04 due to a problem with the PM data sending from the client.

MPP

Done.
104th (13.03.2015);
  • Add a link to the Gdoc for systems not part of the ACCtesting on the MPP website

MPP

 
103th (27.02.2015);
  • Implement changes to SMP before machine checkout

S. Gabourin

 
102th (13.02.2015);
  • Finalize EDMS document for BLM disabling rules for inner triplet monitors

E. B. Holzer

Was presented back in MPP#157, with updated rules for 2018.
101th (30.01.2015);
  • Prepare a template for re-commissioning steps of the SIS to be implemented in ACCTesting

J. Wenninger

 
101th (30.01.2015);
  • Finalize EDMS documents for updated MPS commissioning procedures

MPP

 
99th (28.11.2014);
  • Prepare ECR explaining naming conventions of BLMs in the injection region

E. B. Holzer, B. Dehning

 

98th (31.10.2014)

  • Verify how the BEAM_INFO signal is used for opening/closing logic of the valve

G.Pigny

 

98th (31.10.2014)

  • VSC will provide the information
    how to access the VPI and VGP interlock values to the MKI team.

P.Gomes, J.Uythoven, E.Carlier

Done as confirmed by J.Uythoven by mail on 28.11.2014

97th (24.10.2014)
  • Organise test of the direct dump input

MPP, BLM teams

 

97th (24.10.2014)

  • Organise BLM tests for fast losses (triggering nc circuits) in synergy with FMCM tests

FMCM, BLM teams

 

97th (24.10.2014)

  • Clarify outstanding change request to change the
    orientation of BLM by 90° (requested by T.Baer et al)

C.Bracco

Reported in MPP meeting #98

96th (19.09.2014)
  • Prepare a document with the final implementation of MSS and distribute it for comments.

S. Ravat

C. Martin included these beam permit systems in the checklist for LHC restart. Link to EDMs document describing the layout. DO_14 and DO_15 are used for LHCb and ALICE, ATLAS and CMS use the octo outputs that were already present. 

96th

(19.09.2014)

  • The software development from BI should be followed by an update of the Java application for the AG cleaning and the implementation of the functionality in the SIS.

OP

 

96th

(19.09.2014)

  • An MD for the calibration of the new optical line to be added in the commissioning schedule.

J. Wenninger

 

96th

(19.09.2014)

  • Present a proposal for dump levels to commence operation with in 2015.

J. Uythoven

 

94th

(15.08.2014)

  • Discuss the installation of dBLMs with the different teams involved and propose priorities for installation locations.

D. Wollmann, B. Dehning

 

93rd

(27.06.2014)

  • Introduce orbit bump and/or gas leak scenario for longer RS and compare calculated BLM thresholds with UFO scenario presented.

B. Auchmann, BLMTWG

 

91st (16.05.2014)
  • Review the BCM thresholds outside of stable beams. The higher ALICE can go the lower is the probability to accidentally dump, when colliding the separation bumps. In addition the procedure of setting up collisions in Alice should be improved, most probably by using the luminosity application and by defining an appropriate procedure.

A.Di Mauro, OP

Two operational modes, lower thresholds applied (factor 10) only in stable beams

91st

(16.05.2014)

  • Verify with ALFA team that 45C is really a hard limit for damage of the pots. Discuss and check if the alarms can be implemented into the PM.

S.Wenig, MPP

It was confirmed that overheating was with retracted detectors. The actions taken should improve the situation by a factor of 50. The following decisions were taken

§ Temperature limits will not be included into hardware Injection and/or Beam Permit

§ Temperature limits will not be included in SIS

o It was confirmed that injection is not the critical situation; in fact during the 2 hours between beam dump and injection the temperature goes nearly down to normal values

§ More temperature information will be sent to the machine via DIP; will be included in Timber

§ In case overheating becomes again a problem for detectors despite the actions taken, beam dump will be requested by ATLAS via CCC

o Temperature increase is slow and no urgent actions have to be taken

o This was agreed on during Run-1 with CCC but never used

o Needs to be clarified/officialised with CCC/MPP

 

91st

(16.05.2014)

  • BLM name changes: Gianluca will discuss with Stefano if the changes are finally ok for Collimation, as the change also affects the official name of the BLMs.

COLL, BLM

 

90th (25.04.2014)
  • Verify possibility of compensating a loss of an FGCLite with the OFB or an eventual stratgey to maintain V_REF in the DOC until beams are dumped.

S.Uznanski, MPP

 

90th

(25.04.2014)

  • Establish proposal for commissioing stratgey of CIBUs (which client, what test, when to perform commissioning)

S.Gabourin,C.Martin

 

90th

(25.04.2014)

  • Verify trigger conditions and redundancy of new CIBDS card during reliability run to avoid deadlocks (due to different redundancy implementations in BIS and LBDS)

S.Gabourin, N.Magnin

 

90th

(25.04.2014)

  • Verify SIS logic and signals sent to BIS in CCR and SR3 - Investigate possibility to implement HW links rather than CMW 

S.Gabourin, L.Ponce

 

89th

(04.04.2014)

  • Signal from RF control in LINAC4 is currently not latched - IS team to defined needs and implement changes to protect source against potential of sparking following rapid Beam Permit changes

ABP-IS, RF team

 

89th

(04.04.2014)

  • Propose and document modifiction of Matrix/additional channel to allow for an efficient but safe start-up of the Source

C. Martin, RF team

 

89th

(04.04.2014)

  • Recommission the BIS-RF interface of LINAC4 following modification of RF Card

C. Martin, RF team

 

88th

(21.03.2014)

 

  • Put the ES document for CIBDS card in the circulation.

S. Gabourin

EDMS document in circulation under MPP node

87th (07.03.2014)
  • Check the updated damage limits of the TCTs.

R. Bruce

A paper was published.

87th

(07.03.2014)

 

  • Discuss the SBF limits between different teams. Find out how to do the setup with the proposed SBF limits.

V. Chetvertkova

Discussion being finalized for Evian 2014.

86th

(21.02.2014)

 

  • Update the member-lists and mandate of MPP in preparation for LMC input.

M. Zerlauth

Done and presented in LMC on 14th of May 2014

86th

(21.02.2014)

 

  • Assign a person for making independent checks of commissioning procedures.

MPP

 

86th

(21.02.2014)

 

  • Put together the interlock test procedures in a new EDMS document.

M. Deile

 

85th

(31.01.2014)

 

  • Discuss in MPP the recent case of damage to a bellow during the commissioning of Linac4.

R. Schmidt, B. Puccio, A. Apollonio, D. Wollmann

89th MPP meeting 4th April 2014

85th

(31.01.2014)

 

  • Present update on re-commissioning for the WIC.

M. Zerlauth

 

85th

(31.01.2014)

 

  • Present planned interlock configurations of the PICs to MPP.

I. Romera, M. Zerlauth

Presentation done - final config to be re-iterated with LBOC/ABP? Need to forward information as well to CRYO to assure same masking capabilities!

84th

(13.12.2013)

 

  • Finalize and circulate template of MPS re-commissioning procedures.

M. Zerlauth

Updated Template for MPS procedures

 

84th

(13.12.2013)

 

  • Blind able firmware in the BLM crates: Check with Bernd if firmware for blinable racks can already be implemented for the start-up after LS1 to avoid the need for re-commissioning, if it is implemented later. Follow up with Bernd.

E.B. Holzer, M. Zerlauth

First Draft ECR to be finished by early June.

84th

(13.12.2013)

 

  • BLMs: Define with the equipment teams, which SEMs should be replaced by LICs and where they should remain.

E.B. Holzer, J. Uythoven, S. Redaelli

Feedback from Inj/Dump team sent by Chiara.

IP2, IP6 and IP8 defined. In IR3/7 Stefano tentatively agreed to change some 10 SEMs against LICs for improved monitoring.

82nd

(08.11.2013)

 

  • Warning and dump levels for the BSRA: Jan will write the updated proposal into a note and distribute it..

J. Uythoven

Note remains to be written. Need to talk to OP about required changes to the GUI. We defined the cleaning levels required not to have any quenches, few 1e9 at 7 Tev, so coherent with past experience. However, Jan needs to study what to do with BLM signals etc., as they seem to be incoherent from the past experience (we might just look at lifetime instead).

81st

(25.10.2013)

 

  • Vacuum incident on the MKB dilution kickers: Weigh the recuperated aluminum pieces and estimate the surface size of the missing parts.

A. Vidal

Requested to Vaccuum group but finally not done.

81st

(25.10.2013)

 

  • Define the final levels for warning, dump and no dump.
  • Define the procedure how to switch the abort gap cleaning on and off and how to initiate the beam dump.

J. Wenninger, S. Bart-Petersen, J. Uythoven, S. Mazzoni

Proposal preapred and documented in EDMS document:

https://edms.cern.ch/document/1337184/1 - currently under approval

80th (13.09.2013)
  • Discuss the implementation of the User test modes for Linac4.

C.Martin, Users, MPP

Done.

80th

(13.09.2013)

 

  • Discuss the implementation of the new interlock as  function of separation.

S.Redaelli, A.Masi, J.Wenninger

The case of IP8 needs to be studied. Reserve the 2 telegrams for the moment. A final decision will be taken end of 2013.

79th (19.07.2013)

  • TDI/TCDQ/MSI on BETS: Verify cabling requests for fibres and CIBU connections are taken into account.
  • Check for financing of redundant gap measurement on TDI

ABT and MPE

4 cables asked for BETS’CIBU in UA63, UA67, UA23 and UA87 to the BICs in UA63, UA67, SR2 and SR8 respectively. The cables are foreseen to be installed for the end of October.

78th (05.07.2013)

  • Review commissioning procedures in view of implementation in ACCTEST framework.
  • Define new template layout reflecting upcoming commissioning plans

Equipment teams

MPP

 Type test for exchange of MKD generator being currently implemented. For post LS1 start-up the main tool will remain the sharepoint site.

77th (31.05.2013)

  • Correct inconsistencies in the table and add another column, which describes the proposed changes of BLMs to LICs to deal with injection losses. This table will be circulated for comments

B.Holzer

First table circulated. Discussion with ABT team + Barbara on 04.09.2013. Converged for changes in IR2, Barbara will clean document and proceed as well with IR8.

The final table with the monitor lists for IP2 and IP8 have been distributed by Barbara (IP2_monitors_2014_v0,IP8_monitors_2014_v1)

77th (31.05.2013)

  • Establish test procedure for verification of redundant powering, discuss with hardware commissioning coordinators.

MPP

Type test done on 24th of April, preparation of final procedure ongoing

77th (31.05.2013)

  • Review measured signals in the ICs during injection, extrapolate the signals expected in the LICs (1.1bar) and compare to the factors quoted in Eduardo's presentation.

E.Nebot, W.Bartmann, C.Bracco

 LIC (0.4 bar) 57.7 +-2.0, LIC (1.1 bar) 12.4 +-4.0, FIC assumed identical to LIC 1.1 bar

76th (12.04.2013)

  • Preparation of the ECR for the change of the BLM positions (moving one of the 3 BLMs per quad in the arc to the downstream dipole)

B.Dehning et al

First draft existing, circulated in EDMS (Doc Nr. 1307356)

 

76th (12.04.2013)

  • Trigger a meeting with VAC, RF and OP colleagues to agree on strategy to be adopted for Vast Vacuum valves in IR4.

M.Zerlauth

Presentation of a decision in LMC meeting 166 not to install the vast vacuum valves during LS1.

75th (22.02.2013)

  • Planned displacement of the Access Safety Blocks (ASB) during LS1: Check IR3 radiation map to identify, if there is a need to move the ASBs due to activation to ease maintenance.

S. Hutchins

The RP survey at point 3 has been done (see classification) and shows two hot spots at the TCAPA absorbers:

·  TCAPA6L3B1 : DdD ŕ 500µSv/h au contact - 45µSv/h ŕ 40cm – 22µSv/h ŕ 1m (balisage en local : zone contrôlée séjour limité)

·  TCAPA6R3B2 : DdD ŕ 300µSv/h au contact - 30µSv/h ŕ 40cm – 12µSv/h ŕ 1m (balisage en local : zone contrôlée séjour limité) These are the original positions that were proposed for the ASBs at +/- 177m of IP3. My conclusion would be that it would be unwise to move to these positions in the future as in case of a problem on the ASBs any intervention would incur a higher dose than needed. The current position is satisfactory at present activation levels.

73th (14.12.)

  • Detailed FLUKA studies on beam impact in blade of fast vacuum valves and its impact on downstream accelerator equipment.

FLUKA-team (A. Lechner)

First extrapolation/risk-estimate from wire-scanner quench test done (slides). Numbers quantitatively confirmed by new estimate of C.Garion.

73th (14.12.)

  • Issue with beam transfer between SPS and LHC: The planned changes in the LHC injection scheme, SPS SIS, etc. should be summarized and distributed to the experts of the concerned systems.

J.C. Bau, I. Kozsar

The new LHC central timing that will solve this issue will be deployed the 27th of October.
Timing team will provide a document describing the changes.

72th (07.12.)

  • BLM thresholds: Discuss outlook for 7TeV thresholds in triplets in MPP workshop in March.

BE-BI, MPP

Planned for MPP workshop, 11.-13.03.2013 -> Ongoing Work by BLMTWG

70th (26.10.)
  • Discuss with Stephan Bart Pedersen on post-mortem data of BLMs during injection (diamonds).

D. Wollmann, M. Zerlauth

dBLM data is in PM, follow up on 127th MPP.

70th (26.10.)

  • Hydrodynamic tunnelling experiment in HiRadMat: Presentation of post mortem analysis of the copper target.

J. Blanco, R. Schmidt, F. Burkart, D. Wollmann

The results of the target opening have been presented to in an contributed talk to IPAC13. The presentation and paper are available for download.

FLUKA & BIG2 simulations with the experimental beam parameters are currently on-going and will be finished in late autumn 2013. Their comparison to the measurements will be presented to MPP in due time.

69th (12.10.)

  • Interlocked BPMs in IR6: Follow-up on reflections and post mortem buffer changes.
  • Interlocked BPMs in IR6: Follow-up on reflections and post mortem buffer changes.

E. Calvo, J. Uythoven, M. Zerlauth

Interlocked BPMs will be changed to terminated strip-lines (ECR). Also electronics undergoing review - SW spec to be looked at.

69th (12.10.)

  • Proposed change of BLM positions in the arcs during LS1:
    • Further simulations are needed to optimize the proposed solution of moving one of the 3 BLMs installed at the MQs to the upstream MB.
    • Study the different loss distributions from UFOs.
    • Perform further simulations to study the possible gain from BLMs in the other beam.

BE-BI,

M. Sapinski, A. Lechner

 EDMS doc prepared and approved in LMC (Doc Nr. 1307356)

67th (14.09.)

  • Interlock of LHCb: Check with LHCb if for the interlock of the LHCb magnet the same solution as in the LHC (safety PLC + FPGA) can be used.

M. Zerlauth

Discussion took place with Sylvain Ravat & Colleagues. The MSS for detectors is currently being renovated, based on NI platform. MSS also performs 'quench detection' like functionality which cannot be covered by a PLC.

67th (14.09.)

  • Injection protection: Start brainstorming if and how injection protection can be improved for LS2, especially in the view that there is no passive protection (with collimators, masks etc) in the horizontal plane.

MPP

R. Schmidt

Planned for MPP workshop, 11.-13.03.2013; Postponed for HL-LHC redesign in WP 14

67th (14.09.)

  • Injection protection: For LS1 it is planned to add the TDI jaw position into the MKI BETS, to avoid charging of the MKI, if the TDI is not at the supposed position.

TE-ABT

Discussed and a priori agreed during 79th meeting.

67th (14.09.)

  • Heating od TDI / Scrubbing run 2012: Before the scrubbing run a sequence step needs to be created. Markus proposes to make a special sequence for the scrubbing run, which reduces the likelihood of moving the TDI out without putting the MKI into STANDBY.

V. Kain

C. Bracco

Was implemented and used as described for 25ns scrubbing end 2012.

66th (31.08.)

  • Linac4 failure catalogue: Meet with the different experts to get the missing information.

A. Apollonio

EDMS doc Nr. 1016233 + commissioning procedure

66th (31.08.)

  • Linac4 BIS: Finalize BIS specification.

B. Mikulec, B. Puccio

EDMS doc Nr. 1016233

66th (31.08.)

  • Linac4 BIS: Discussion strategy for mask-able and un-mask-able inputs to the BIC.

B. Mikulec, B. Puccio, R. Schmidt, M. Zerlauth

EDMS doc Nr. 1016233

65th (17.08.)

  • CMS solenoid: Add fast power abort of CMS solenoid to CMS beam permit, to avoid its recently observed influences on the beam (beams were dumped due to high losses in IR7). Discuss with N. Bacchetta in September.
    • Discuss with N. Bacchetta in September.
    • Present possible solutions to MPP.
  • Idem for recent rip of LHCb power supply trip. Sylvain Ravat to investigate possibilities to decrease the filter time + relay activation time of the MSS (10+15ms)

 

N. Bacchetta, S. Ravat, D. Wollmann, M. Zerlauth

CMS features an MSS (Magnet Safety System) + CIBU, the latter is however currently not connected. A HW Signal (indicating a Fast Discharge or Slow Discharge event in the solenoid) exists already. Everyone agreed however that only a Fast Discharge should be considered to eventually dump the beams. All 4 LHC experiments used the same MSS, due to obsolete components (Fuse Programmable ALTERA FPGA), a new design is already undergoing for LS1 (based on National Instruments Rio card). Upon this occasion a HW Signal for Fast Discharges will be provided, if and when this Signal will then be connected to the CIBU can be decided as a Function of other mitigation measures available after LS1 (slow OFB during SB, SIS on slow changes,). 
-) The LHCb trip was correctly reported by the MSS but has an unusual delay of 25 ms after the initial failure of the power converter (no big orbit changes nor losses observed yet). This was traced back to input filters and additional safety relays at the output of the MSS. Due to the on-going re-design of the MSS (see above), the involved colleagues will study the possibility of using opto-couplers instead of slow relays + reduce the filtering needs (by replacing voltage signals with current loops for less EMC sensitivity). It is expected to bring down the additional delay to ~5-10 ms.

65th (17.08.)

  • LS1 work on the upgrade of the UPS power distribution of the LBDS:
    • TE-MPE shall provide the redundant link from the BIS to the re-triggering lines of the LBDS (allowing to trigger a delayed asynchronous dump in case the LBDS doesn't cause a synchronous one within 250ms).
    • The LBDS UPS powering upgrade should be re-addressed in MPP next year to report on the status of the mitigations.

TE-MPE

MPP

See MPP meeting 77

64th (13.07.)

  • PSpice simulation of the converter trip in RTQX2.L2:
    • Correlate the current change and the orbit change, with the help of the post mortem data. (MZ, ER)
    • Discuss possibilities to improve the circuit current supervision in the triplet, to ease the understanding of these types of events.
    • Discussed presented results with Hugues and Valerie.
    • Decrease the validation time of an I_ERR deviation (currently 2 seconds). To be done during LS1.

M. Zerlauth,
E. Ravaioli,
V. Akpata-Buchon

Simulated current changes seem consistent with orbit recordings in PM.

Possibilities for adding additional LEMs in the free-wheeling paths of the triplet circuit have been discussed with EPC. An Ethernet based data acquisition system (like for the SVC) might be adapted for this purpose. To be followed by EPC.

 

ECR written for changed over-current thresholds: https://edms.cern.ch/document/1227985/1

 

64th (13.07.)

  • Increase BLM thresholds in MQYs (Q4L6, Q4R6) and MQWs (Q5R7, Q5L7):
    • Estimate (and monitor) the increase in integrated dose on the MQWs due to the change in the BLM thresholds, to ensure that the lifetime of the magnets is not reduced by this change. Discuss with Ezio.
    • Ask Ezio to give a presentation on the expected lifetime of the warm magnets in IR7 due to constant beam losses. What are the worries, what are the results of the resent radiation dose measurements

MPP, E. Todesco

Francesco Cerutti is currently working on the interpretation of the data measured with ~50 sensors in the LHC. The results will be presented in the Collimation Working Group (CWG) on the 3rd of September at 16h (874-1-11). MPP members concerned by this topic are of course encouraged to join the meeting.

Improved regulation library to be deployed ahead of the LHC re-start (if FGC2 memory allows). To be confirmed by Q.King.

63rd (22.06.)

  • BPM sensitivity (IR6): Eva will replace all attenuators of BPMSA.A4R6.B2, BPMSA.B4R6.B2 and BPMSB.A4R6.B1 and re-validate the changes with a scraping test. Eva will install the new firmware during the BI MD to perform first tests, but then roll back to the current version during the TS. The firmware will then be installed for monitoring purpose in a parallel crate without being connected to the BIS (and possibly deployed in the operational systems when its functionality has been verified).

E. Calvaro

Item covered by more recent discussion on overall strategy and design spec (see actions from MPP # 69)

63rd (22.06.)

  • FLUKA simulations on beam impact during asynch. dump failure: Mario and Anton will cross-check the input parameters (number of impacting bunches, impact parameters, beam sizes,…) and compare the results of the current simulation with previous estimates for beam impacts on the XRPs.

A. Lechner,
M. Deile

The different results can be mainly explained by different beam sizes and energies used in the two cases (4TeV beta*=0.6m versus 7TeV and beta*=1540m). See also Anton's slide from the AOB of the 65th MPP (17.08.2012).

63rd (22.06.)

  • UPS replacement project:
    • EN-EL and TE-MPE to study all UPS failure modes and the impact on the PIC logic to be implemented. The new logic must be tested while exercising all possible failure cases together with the MPE colleagues on a test-bed before deployment in the LHC.
    • TE-MPE has to provide the requirements for powering after LS1 (specially whether the use of F4 line remains restricted to QPS clients or whether one may relax this constraint).

M. Zerlauth,

TE-MPE

  • TE-MPE has given the authorization to use the F4 as well for redundant powering of other machine protection systems as of LS1 (e.g. LBDS, BLMs,..).
  • The required connections should however be announced to TE-MPE in order to verify the compatibility with the existing QPS system.

63rd (22.06.)

  • Analysis on power converter trip of RTQX2.L2 on 15.06.2012:
    • Arjan and Jorg will try to estimate the current changes actually seen by the magnets, once through simulations and secondly deriving it from the orbit excursions observed.
    • Hugues will reduce the over current protection thresholds LIMITS.I.POS and the allowed current error I_ERR to more adequate values for 4TeV operation. An ECR is required.

A. Verweij,
J. Wenninger,
H. Thiesen

Results of simulations were presented at the 64th MPP. For details see corresponding presentations and minutes.

62nd (01.06.)

  • FLUKA simulations on beam impact during asynch. dump failure:
    • Simulate the case with a direct impact into the stainless steel foil (Anton). Daniel will provide the beam parameters for the case.
    • Due to a mistake in the simulation the deposited energy was underestimated by a factor ~2. Anton corrected this and will present the corrected results in the upcoming MPP.

A. Lechner, D. Wollmann

See presentations and minutes from the 63rd MPP.

62nd (01.06.)

  • Increase BLM thresholds in IR7:
    • Eduardo will check if the reduction of the BLM threshold (TCPs, TCSGs) in the shorter running sums could cause a problem and report to MPP.
    • Belen and Eduardo will prepare a proposal to perform a comparable increase of BLM thresholds in IR3

E. Nebot, B. Salvachua

Proposed BLM threshold changes (IR3) were presented at the 64th MPP. For details see corresponding presentations and minutes.

62nd (01.06.)

  • RP setup and operation: Mario, Daniel and Stefano will meet to create an EDMS document for the RP setup and operation in the different beam modes.

M. Deile, S. Redaelli, D. Wollmann

See EDMS document 1225356

62nd (01.06.)

  • MD classification: Markus and Daniel will send an email to the colleagues responsible for MDs classified as C or D to remind them to provide a respective MD document.

D. Wollmann, M. Zerlauth

 

61st (11.05.)

  • SMP cross checker:
    • In the lab all tests were successfully performed. It would be good to also perform a test in the LHC. This could be done by sending a fake timing event.
    • Include the qualification of the Safe Beam flag in the cross checker (for after LS1).

TE-MPE-MS

S. Gabourin, B. Todd

  • SMP cross-checked tests performed in TS3 (S. Gabourin - see details)
  • B. Todd: Having studied the various options for a software based cross-check of the output of the CISV, the benefit to the machine in terms of increased safety is not balanced by the effort involved in making the cross-check, and the increased risk of unavailability. It’s a very difficult thing to do reliably… We decided not to make the cross-check.

60th (27.04.)

  • BLM HV issue in IR7:
    • Tests showed that the ionization chambers still work down to a HV of 200V. Saturation due to space charge effects might start earlier. Barbara will distribute the documentation of these tests.
    • Functionality tests with a lower HV are foreseen with beam after TS1. The tests will be performed left and right of IR7, i.e. with B1 and B2. The data recorded with the lower HV need to be compared to data without changes, to guarantee the protection functionality of the BLM system.

E.B. Holzer,

BE-BI-BL

Mitigation strategy defined and accepted in MPP # 73

60th (27.04.)

  • Performance of LIC detectors in the LHC and planned replacement of SEMs by LICs for certain areas: Follow-up by MPP and re-address the applicability of LICs in certain areas of the LHC in a future MPP meeting.

MPP

BE-BI-BL

E. Nebot del Busto: Regarding the LICs installed at the MSI locations, the signals observed by this monitors are quite consistent with
those observed by ICs in near locations (just ~60 cm
from the closest LIC). See a few slides attached that try to clarify the location of the detectors and the signals observed during several injections (on the 6th of June). From this we conclude that these detectors are working as expected. However, the losses observed at this location are relatively low. The performance of the LICs will be tested in a HiRadMat experiment (27.-31.08.).

Reason for initial problems now understood and presented in MPP # 70

60th (27.04.)

  • Interlocking of the TDI position during the different modes of operation. Do we need to have other / more interlocks to ensure the correct position of the TDI?
    • The parking position of the TDI during physics was changed during 2011 from 20 to 55 mm (gap). A comparison of the heating / deformation for these two gaps may help to understand the sources of the heating. Jan will look into this

J. Uythoven

J. Uythoven: We started the year with a TDI retraction during physics of around 35 mm instead of fully out at 55 mm. This was an operational mistake. In May I concluded that we are heating a lot less at 55 mm during parking (see plots).

The TDIs are now operated at their fully out position of 55mm during physics.

 


Last updated 28-Feb-2018 17:47